OpenStudy (anonymous):

Find the value of x so that the line passing through (x, 10) and (-4, 8) has a slope of

OpenStudy (anonymous):

how do i find x?

OpenStudy (cwrw238):

as a slope of what?

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

\[\frac{10-8}{x+4}=\frac{2}{3}\] solve for \(x\)

OpenStudy (anonymous):

you can just about do it in your head since \(10-8=2\) you have \[\frac{2}{x+4}=\frac{2}{3}\] so \(x+4=3\)